OpenStudy (anonymous):

What does (coefficient of x/2)^2 = last term mean? For ex. (x-1)^2 = (x-1)(x-1) = x^2 -2x + 1 (right side is (coefficient of x/2)^2 = last term.

OpenStudy (anonymous):

Additionally, (x-7)^2 = (x-7)(x-7) = x^2 - 14x + 49 is true that (coefficient of x/2)^2 = last term.

OpenStudy (anonymous):

i cannot make heads or tails of this question

OpenStudy (anonymous):

ooooooooooooh yes i can i think if you have for example \[x^2-14x+49\] then the "coefficent of the x term" is \[-14\] and \[-\frac{14}{2}=-7\] so you get \[(x-7)^2=x^2-14x+49\]

OpenStudy (anonymous):

ohhhh. thank you!